# Bulk plasmons in elemental metals

### v1

@@ III Results
Table 1: List of properties of the 25 elemental metals studied in this work. The experimental lattice parameters and the corresponding unit cell volume are taken from the compilation made in Ref. [70].
III.1 Intra-band v.s. inter-band contributions
@@
Figure 1: Spectral contributions of the intra-band (dashed orange line), the inter-band (dash-dotted blue line), and the plasma (solid green line) frequency as a function of energy, evaluated with Eqs. (7), (8), and (9), respectively. The total area under the plasma curve is filled with each corresponding contribution.
Table 1 reports the respective intra- and inter-band frequencies, and the percentage of intra-band contributions to the plasma frequency for the 25 studied metals. While the intra-band frequencies of all the metals range from 3.74 eV for Ti to $11.84$ eV for Al, the inter-band frequencies have a far wider range, from only $1.42$ eV for Na to $50.89$ eV for W. This huge span is due to inter-band contributions of d- and f-orbitals, thus intra-band contributions are relatively more relevant for metals with only $s$ and $p$ electrons. The plasmon of Na is dominated by intra-band contributions (94.0 %), followed by Li (66.5 %), and Al (62.2 %). Next in the list are Mg, with less than half (47.1 %), and K (13.4 %). The percentage for the rest of the metals is much lower.
@@ III.2 Spectra in the optical limit
Figure 2: Comparison of the computed RPA loss function of V (a), Cu (b), and Zn (c) in the optical limit, with experimental EELS data from Ref. [54], REELS from Ref. [26] and theoretical IPA calculations from Ref. [26].
In Fig. 2 we compare the computed RPA loss function of V, Cu, and Zn in the optical limit, with EELS measurements from the EELS Atlas [54] and REELS data and IPA from Ref. [26]. These are examples of materials with complex screening effects resulting in several peaks in their response function, that require a description beyond a single pole PPA model [57, 4, 31].
The raw EELS data from Ref. [54] includes a background intensity from the zero-loss peak, which we removed using a standard power law model [58, 11], while the remaining intensity was normalized according to the maximum RPA intensity. These measurements have a lower resolution and a larger instrumental broadening than those of Ref. [26], however, for most cases, the energy range is larger, providing information on the tail of the spectra. The experimental loss function of Ref. [26] was obtained by performing Kramers-Kronig transformations [26, 11] on the raw REELS data. The results in the tail region are therefore affected by the finite energy interval in the frequency integral of such transformations, and in several cases, like Cu, are less accurate than the direct EELS measurements.
In all cases, both our RPA calculations and the IPA results from Ref. [26] are able to accurately describe the main features in the experiments. The inclusion of local field effects with RPA improves the description of the intensity with respect to IPA. Refining the DFT starting point, e.g., with DFT+U or hybrid functionals, may improve the description of the finer features in the loss function of materials with $d$-states and strong spin-orbit interactions [78, 79, 80, 55]. Such refinements are beyond the scope of this work.
@@ III.3 Deviation from the free-electron gas and PPA
Figure 4: (a) Relation between the effective number of electrons, $Z_{\mathrm{eff}}$, and the number of valence electrons, $Z_{\mathrm{val}}$, listed in Table 1 for the set of studied metals. $Z_{\mathrm{eff}}$ is computed from the energy position of the main MPA pole, with Eq. (18). The black dashed line represents the identity relation. (b) Relation between the spectral weight, $\mathrm{Re}[2R]$, and the energy position, $\mathrm{Re}[\Omega]$, of the main MPA pole of all the studied metals. The black dashed line corresponding to $\mathrm{Re}[2R]=\mathrm{Re}[\Omega]$, represents the ideal PPA model where all the spectral weight is concentrated in a single pole. The presence of other poles with significant spectral weights lowers the position with respect to this PPA line. The gray dashed line represents a pole with half the spectral weight.
The free electron gas model has often been used to describe simple metals like Al and Na [1, 2, 81], considering an electronic density given by the number of their valence electrons. Such an approach can be used for other metals, but it is not always clear which electronic density to consider. With MPA in place, we can define an effective number of electrons from the expression of the classical plasmon energy in Eq. (11):
@@
Fig. 4(a) shows the relation between the computed $Z_{\mathrm{eff}}$ and the number of valence electrons, $Z_{\mathrm{val}}$, for the set of 25 studied metals. The values are listed in the last column of Table 1. For some of the cases $Z_{\mathrm{eff}}$ is close to the number of electrons in the outer orbitals. For example, the value of $0.9$, $2.1$, and $10.17$ are very close to $1$, 22, and $10$, corresponding to the number of electrons in the 3s1, 4s2, and 4d10 orbitals of Na, Ca, and Pd, respectively. However, in several cases like Co, Ag, and Ta, the mixed nature of the different orbital contributions to the plasmon energy and thus $Z_{\mathrm{eff}}$, prevents us from establishing any simple association.
On the ab initio side, the single pole PPA model, has been extensively used in $GW$ and similar approaches beyond DFT, even for systems with several prominent poles in their response functions. The MPA representation captures such a multipole behavior, and the spectral weight of its dominant pole provides a quantitative measure of how far is the PPA description.
Fig. 4 shows the spectral weight and the energy position of the main MPA pole in the response function of the 25 studied metals. According to Eq. (10), in PPA the spectral weight should equal the plasmon energy, represented by the black dashed line. The gray dashed line indicates half of the total spectral, $\mathrm{Re}[2R]=\mathrm{Re}[\Omega]/2$. The results show that Na, Mg, Al, K, and Sn are the systems well described by PPA, and the plasmon of Al has the largest spectral weight of all the metals. For Be, Li, Ca, W, Ta, Pb, and Tl the main pole has nearly half the total spectral weight, while for the remaining metals has even smaller weights. Ag has the smallest spectral weight and one of the largest deviations from PPA.
@@ III.4 Spectral $Y({\mathbf{q}},\omega)$ band structures
Fig. 5 shows the dispersion of the main excitations in $Y({\mathbf{q}},\omega)$, along two high-symmetry $\mathbf{q}$}-lines of bulk Li (a), Na (b), Al (c), K (d), and Ca (e). The selected $\mathbf{q}$}-paths correspond to $H\Gamma N$ and $X\Gamma N$ for the bcc and fcc Bravais lattices, respectively. Such simple elemental metals show a parabolic-like dispersion of the main plasmon, with an effective mass that increases with the plasmon energy. Secondary excitations are more visible for heavier elements, showing a much flatter dispersion than that of the main plasmon.
@@
Fig. 6 shows analogous plots for V (a), Cr (b), Fe (c), Ni (d), and Cu (e), while Fig. 7 shows plots along the $HA\Gamma M|K\Gamma$ path for metals with hexagonal symmetry: Be (a), Ti (b), and Co (c). The complexity of the excitation landscape of the heavier metals increases very quickly with respect to the cubic elements in Fig. 5 and the light hexagonal ones (Be and Mg). The main plasmon is split into a multiple-peak structure with mostly flat dispersions, at variance with what is expected from the free electron gas model with an equivalent electronic density. There are also many more secondary excitations with a significant spectral weight, as discussed in the previous section. The dispersion of such excitations is also more complex, and gives rise to diverse features in the spectral band structure, as discussed in the following.
The plasmonic dispersions are in general anisotropic and may have discontinuities due to anisotropies in the electronic band structure [64, 82, 67, 83]. Examples of such discontinuities are visible in Fig. 7 at the $\Gamma$ point of the $A\Gamma M$ path, for instance, around $30$ eV for Be, 4545-$55$ eV for Ti, and 2020-$35$ eV for Co. In the latest case, the discontinuity extends to the region of the plasmon, similar to other hexagonal systems with $d$-electrons showing strong anisotropy signatures, such as ZnO [55]. Other spectral features are present in many of the studied metals, such as non-parabolic energy and intensity dispersions [84, 83], including negative dispersions and indirect excitations, i.e., with a finite intensity at finite $\mathbf{q}$}, that vanishes at ${\mathbf{q}}=\Gamma$. There is also quasi-particle overlapping leading to band crossings and anti-crossings, as shown in Fig. 6 in the zoomed region around $10$ eV.
Individual plots analogous to the ones in Figs. 5, 6, and 7, for all the 25 metals studied here, can be found in Ref. [85]. An effective analytical representation of the spectral $Y({\mathbf{q}},\omega)$ band structure and the interpretation of its features is presented in the next section.
@@ III.5 Spectral properties within MPA($\mathbf{q}$})
Figure 8: Comparison of the spectral $Y({\mathbf{q}},\omega)$ band structure of V in the $\Gamma N$ direction, generated from the numerical RPA data with its corresponding MPA model with $n_{Y}=13$ poles.
Many of the studied metals exhibit spectral $Y({\mathbf{q}},\omega)$ band structures with a high degree of complexity, as illustrated in the previous section. Still, one can track the dispersion of the most prominent peaks, which varies smoothly with $\mathbf{q}$}. This suggests that such a complex spectrum can be modeled with a small set of plasmonic excitations with a simple $\mathbf{q}$} dependence.
To test this hypothesis, we have fitted MPA($\mathbf{q}$}) models, with a number of poles of $n_{Y}=2$ to $n_{Y}=15$, to the numerical $Y({\mathbf{q}},\omega)$ data of the 25 metals studied in this work. The fit is performed along a selected high symmetry $\mathbf{q}$}-line, $\Gamma N$ for cubic systems and $\Gamma M$ for hexagonal. Analogous fits can be made along the other $\mathbf{q}$}-paths. The results can be found in Ref. [85]. In all the cases, the resulting $Y^{\mathrm{MPA}}({\mathbf{q}},\omega)$ reproduces very accurately the numerical data, even for systems with complex pole structures.
@@
Fig. 8 shows the case of V, which is fitted with the same number of $n_{Y}=13$ poles than in Fig. 3. The numerical spectral band structure in Fig. 8(a) is accurately reproduced with MPA($\mathbf{q}$}) in Fig. 8(b), while Fig. 8(c) includes dashed black lines indicating the energy position of the poles, $\mathrm{Re}[\Omega_{p}({\mathbf{q}})]$. Analogous plots for the rest of the 25 metals are provided in Ref. [85]. The analytical MPA($\mathbf{q}$}) representation also offers the possibility to disentangle overlapping poles, which in many cases is not trivial. In the following, we use Ca and Ni to exemplify such untangling.
Fig. 9(a) shows the total spectral band structure of Ca, reconstructed with MPA($\mathbf{q}$}) with 6 poles, analogous to the case of V in Fig. 9(c). In panels (b-g) we show the contribution of each individual pole to the sum in Eq. (14). The poles are labeled according to their energy position at ${\mathbf{q}}=0$. The resulting MPA($\mathbf{q}$}) poles of Ca present a sufficiently small broadening and smoothly varying intensity to allow identifying them as well defined bands along the whole $\mathbf{q}$} range. The spectrum shows two main poles in line with Ref. [57]. For most values of $\mathbf{q}$}, the most intense one is $p=5$, while $p=4$ dominates around ${\mathbf{q}}=N$. Both poles present a non-monotonic spectral weight with a maximum at finite $\mathbf{q}$}. The three upper poles ($p=4,5,6$) cross at a $\mathbf{q}$}-point slightly before half the $\Gamma N$ distance. The maximum total intensity is located around this region, as a consequence of a constructive pole superposition: $\mathrm{Re}[R_{p}({\mathbf{q}})]>-\mathrm{Im}[R_{p}({\mathbf{q}})]>0$, reflecting the constraints in Eq. (16) for the pole positions $\Omega_{p}$.
@@
In Fig. 10, we show analogous MPA($\mathbf{q}$}) spectral band structures for Ni, in this case modeled with 13 poles. As for Ca, the most intense poles are well defined and superimpose constructively with the intensity spreading in a larger energy range. However, many poles, especially in the low-energy region (p=2p=2-66) around $10$ eV shown in Fig. 6, exhibit residues with an imaginary part larger than its real part and a sign not always complying with the time ordering. Imposing such constraints on the residues $R_{p}$, using a relation analogous to the one for the poles in Eq. (16), results in a reduction of the accuracy of the MPA($\mathbf{q}$}) fit with respect to the numerical data. The need for a fit with free residues, which allows for incoherent pole interference, evidences the complexity of plasmonic quasi-particle formation in non-homogeneous electronic materials.

### v2

@@ III Results
Table 1: List of properties of the 26 elemental metals studied in this work. The experimental lattice parameters and the corresponding unit cell volume are taken from the compilation made in Ref. [2].
III.1 Intra-band vs inter-band contributions
@@
Figure 1: Spectral contributions of the intra-band (dashed orange line), the inter-band (dash-dotted blue line), and the plasma (solid green line) frequency as a function of energy, evaluated as running integrals of Eqs. (7), (8), and (9), respectively. The total area under the plasma curve is filled with each corresponding contribution.
Table 1 reports the respective intra- and inter-band frequencies, and the percentage of intra-band contributions to the plasma frequency for the 26 studied metals. While the intra-band frequencies of all the metals range from 3.74 eV for Ti to $11.84$ eV for Al, the inter-band frequencies have a far wider range, from only $1.42$ eV for Na to $50.89$ eV for W. This huge span is due to inter-band contributions of d- and f-orbitals, thus intra-band contributions are relatively more relevant for metals with only $s$ and $p$ electrons. The plasmon of Na is dominated by intra-band contributions (94.0 %), followed by Li (66.5 %), and Al (62.2 %). Next in the list are Mg, with less than half (47.1 %), and K (13.4 %). The percentage for the rest of the metals is much lower.
@@ III.2 Spectra in the optical limit
Figure 2: Comparison of the computed RPA loss function of V (a), Cu (b), and Zn (c) in the optical limit, with experimental EELS data from Ref. [1], REELS from Ref. [84], and theoretical IPA calculations from Ref. [84].
In Fig. 2 we compare the computed RPA loss function of V, Cu, and Zn in the optical limit, with EELS measurements from the EELS Atlas [1] and REELS data and IPA of Ref. [84]. These are examples of materials with complex screening effects resulting in several peaks in their response function, that require a description beyond a single pole PPA model [58, 60, 47].
As mentioned in Sec. II.3, the raw EELS data provided in Ref. [1] includes a background intensity from the zero-loss peak, was removed using a standard power law model [79, 18], while the remaining intensity was normalized according to the maximum RPA intensity. These measurements have a lower resolution and a larger instrumental broadening than those of Ref. [84]; however, for most cases, the energy range is larger, providing information on the tail of the spectra. The experimental loss function of Ref. [84] was obtained by performing Kramers-Kronig transformations [84, 18] on the raw REELS data. The results in the tail region are therefore affected by the finite energy interval in the frequency integral of such transformations, and in several cases, like Cu, are less accurate than the direct EELS measurements.
In all cases, both our RPA calculations and the IPA results of Ref. [84] are able to accurately describe the main features of the experimental data. The inclusion of local field effects with RPA improves the description of the intensity with respect to IPA. Refining the DFT starting point, e.g., with DFT+U or hybrid functionals, may improve the description of the finer features in the loss function of materials with $d$-states and strong spin-orbit interactions [87, 67, 65, 46]. Such refinements are beyond the scope of this work.
@@ III.3 Deviation from the free-electron gas and PPA
Figure 4: (a) Relation between the effective number of electrons, $Z_{\mathrm{eff}}$, and the number of valence electrons, $Z_{\mathrm{val}}$, listed in Table 1 for the set of studied metals. $Z_{\mathrm{eff}}$ is computed from the energy position of the main MPA pole, with Eq. (18). The black dashed line represents the identity $Z_{\mathrm{eff}}=Z_{\mathrm{val}}$. (b) Relation between the spectral weight, $\mathrm{Re}[2R]$, and the energy position, $\mathrm{Re}[\Omega]$, of the main MPA pole of all the studied metals. The black dashed line corresponding to $\mathrm{Re}[2R]=\mathrm{Re}[\Omega]$, represents the ideal PPA model where all the spectral weight is concentrated in a single pole. The presence of other poles with significant spectral weights lowers the position with respect to this PPA line. The gray dashed line represents a pole with half the spectral weight.
The free-electron gas model has often been used to describe simple metals like Al and Na [7, 39, 85], considering an electronic density given by the number of their valence electrons. Such an approach can be used for other metals, but it is not always clear which electronic density to consider. With MPA in place, we can define an effective number of electrons from the expression of the classical plasmon energy in Eq. (11):
@@
Fig. 4(a) shows the relation between the computed $Z_{\mathrm{eff}}$ and the number of valence electrons, $Z_{\mathrm{val}}$, for the set of 26 studied metals. The values are listed in the last column of Table 1. For some of the cases $Z_{\mathrm{eff}}$ is close to the number of electrons in the outer orbitals. For example, the value of $0.9$, $2.1$, and $10.17$ are very close to $1$, 22, and $10$, corresponding to the number of electrons in the 3s1, 4s2, and 4d10 orbitals of Na, Ca, and Pd, respectively. However, in several cases like Co, Ag, and Ta, the mixed nature of the different orbital contributions to the plasmon energy and thus $Z_{\mathrm{eff}}$, prevents us from establishing any simple association.
On the ab initio side, the (single-pole) PPA model, has been extensively used in $GW$ and similar approaches beyond DFT, even for systems with several prominent poles in their response functions. The MPA representation captures such a multipole behavior, and the spectral weight of its dominant pole provides a quantitative measure of the unsuitability of a (single-pole) PPA description.
Fig. 4 shows the spectral weight and the energy position of the main MPA pole in the response function of the 26 studied metals. According to Eq. (10), in PPA the spectral weight should equal the plasmon energy, which is represented by the black dashed line. As a guide to eye, the gray dashed line indicates half of the total spectral, $\mathrm{Re}[2R]=\mathrm{Re}[\Omega]/2$. The results show that Na, Mg, Al, K, and Sn are the systems well described by a (single-pole) PPA model, and the plasmon of Al has the largest spectral weight of all the metals. For Li, Be, Ca, W, Os, Ta, Tl, and Pb the main pole carries nearly half the total spectral weight, whereas for the remaining metals it contributes even less, with Ag exhibiting the smallest spectral weight and representing one of the systems furthest from a (single-pole) PPA picture.
@@ III.4 Spectral $Y({\mathbf{q}},\omega)$ band structures
Fig. 5 shows the dispersion of the main excitations in $Y({\mathbf{q}},\omega)$ along two high-symmetry $\mathbf{q}$}-lines of bulk Li (a), Na (b), Al (c), K (d), and Ca (e). The selected $\mathbf{q}$}-paths correspond to $H\Gamma N$ and $X\Gamma N$ for the bcc and fcc Bravais lattices, respectively. Such simple elemental metals show a parabolic-like dispersion of the main plasmon, with an effective mass that increases with the plasmon energy. Secondary excitations are more visible for heavier elements, showing a much flatter dispersion than that of the main plasmon.
@@
Fig. 6 shows analogous plots for V (a), Cr (b), Fe (c), Ni (d), and Cu (e), while Fig. 7 shows plots along the $HA\Gamma M|K\Gamma$ path for metals with hexagonal symmetry: Be (a), Ti (b), and Co (c). The complexity of the excitation landscape of these metals increases very quickly with the atomic number, particularly when valence $d$-electrons are present, contrasting with the light cubic elements in Fig. 5 and the light hexagonal ones (Be and Mg). In these cases, the main plasmon is split into a multiple-peak structure with mostly flat dispersions, in contrast to what would be expected from the free-electron gas model with corresponding electronic densities. In addition, there are many more secondary excitations with a significant spectral weight. The dispersion of such excitations is also more complex, and gives rise to diverse features in the spectral band structure.
Complex features in the spectra include anisotropic plasmonic dispersions and discontinuities that arise due to anisotropies in the underlying electronic band structure [62, 80, 77, 15]. Examples of such discontinuities are discernible in Fig. 7, such as for Be around $30$ eV at the $\Gamma$ point along the $A\Gamma M$ path, and correspondingly at 4545-$55$ eV for Ti, and 2020-$35$ eV for Co. For the case of Co, the discontinuity extends to the region of the plasmon, similar to other hexagonal systems with $d$-electrons showing strong anisotropy signatures, such as ZnO [46]. Other spectral features, such as non-parabolic energy and intensity dispersions [35, 15], including negative dispersions and indirect excitations, i.e., with a finite intensity at finite $\mathbf{q}$}, that vanishes at ${\mathbf{q}}=\Gamma$, can be seen in many of the studied metals. There is also quasi-particle overlapping leading to band crossings and anti-crossings, as shown in Fig. 6 in the zoomed region around $10$ eV.
@@ III.5 Spectral properties within MPA($\mathbf{q}$})
Figure 8: Comparison of the spectral $Y({\mathbf{q}},\omega)$ band structure of V in the $\Gamma N$ direction, generated from the numerical RPA data with its corresponding MPA($\mathbf{q}$}) model with $n_{Y}=13$ poles.
Spectral $Y({\mathbf{q}},\omega)$ band structures for each of the studied 26 metals are provided in Ref. [3]. Many of them exhibit complex features, analogous to those in Fig. 7. Nonetheless, at least for the most prominent peaks, the dispersion with $\mathbf{q}$} is smooth enough to be easily followed. This property suggests that $Y({\mathbf{q}},\omega)$ can be modeled with a small set of plasmonic excitations with simple $\mathbf{q}$} dependence. To test this hypothesis, we have fitted an MPA($\mathbf{q}$}) model to the numerical $Y({\mathbf{q}},\omega)$ data of each of the 26 metals studied, using a number of poles that varies from $n_{Y}=2$ to $n_{Y}=15$, depending on the complexity of the spectra. The fit is performed along a selected high symmetry $\mathbf{q}$}-line, $\Gamma N$ for cubic and $\Gamma M$ for hexagonal systems. Analogous fits can be made along the other $\mathbf{q}$}-paths. In all the cases, the resulting $Y^{\mathrm{MPA}}({\mathbf{q}},\omega)$ reproduces very accurately the numerical data, as shown in Ref. [3] and in the following for the specific examples of V, Ca, and Ni.
@@
Figure 10: Relation between the real and the imaginary parts of the poles $\Omega_{p}(q)$ (a) and the residues $2R_{p}(q)$ (b) of the MPA($\mathbf{q}$}) representation of Ca, in the complex frequency plane.
Figure 8 illustrates the case of V. The numerical $Y({\mathbf{q}},\omega)$ spectral band structure, shown in Fig. 8(a), is accurately reproduced with a fitted MPA($\mathbf{q}$}) model with $n_{Y}=13$ poles, the same number used in Fig. 3, as shown here in Fig. 8(b). To facilitate the interpretation, Fig. 8(c) plots the same MPA($\mathbf{q}$}) model with dashed black lines indicating the energy position of the poles, $\mathrm{Re}[\Omega_{p}({\mathbf{q}})]$.
The analytical MPA($\mathbf{q}$}) representation also provide means to disentangle overlapping poles, which is highly nontrivial. Figure 9(a) shows the total spectral band structure of Ca, reconstructed with MPA($\mathbf{q}$}) with 6 poles. Figures 9(b)-(g) shows the individual contribution of each pole to the sum in Eq. (14), labeled according to their energy position at ${\mathbf{q}}=0$. The MPA($\mathbf{q}$}) poles of Ca have sufficiently small broadening and smoothly varying intensity to allow identifying them as well-defined spectral bands along the whole $\mathbf{q}$} range. The spectrum shows two main poles in line with Ref. [58]. For most values of $\mathbf{q}$}, the most intense pole is $p=5$, while $p=4$ dominates around the ${\mathbf{q}}=N$ high-symmetry point. Both poles present a non-monotonic spectral weight with a maximum at finite $\mathbf{q}$}. The three upper poles ($p=4,5,6$) cross at a $\mathbf{q}$}-point slightly before half the $\Gamma N$ distance. The maximum total intensity is located around this region, as a consequence of a constructive pole superposition.
In Fig. 10 we show the position of the poles $\Omega_{p}(q)$ (a) and the residues $2R_{p}(q)$ (b) in the complex frequency plane. The poles satisfy the constraint of Eq. (16), which is indicated by the gray area in Fig. 10(a), as imposed in the fitting. Even if no constraints are imposed on the residues, they locate in the gray area of Fig. 10(b), representing a similar condition, $\mathrm{Re}[R_{p}(q)]>-\mathrm{Im}[R_{p}(q)]>0$, except for $p=4$, which is slightly outside in the limit $q\to 0$ [See the parameters of the MPA($\mathbf{q}$}) representation in Table 2]. This result indicates the constructive character of the pole superposition of Ca.
@@
$\Omega_{p}$~(eV) $\Omega^{\prime}_{p}$ $\Omega^{\prime\prime}_{p}$ $\Omega^{\prime\prime\prime}_{p}$ $R_{p}$~(eV) $R^{\prime}_{p}$ $R^{\prime\prime}_{p}$ $R^{\prime\prime\prime}_{p}$
Ca ($n_{Y}=6$)
2.161.71i2.16-1.71i 0.480.14i0.48-0.14i 0.50+0.23i-0.50+0.23i 0.40+0.12i-0.40+0.12i 0.100.00i0.10-0.00i 0.960.10i0.96-0.10i $0.58+0.03i$ $0.24+0.12i$
3.750.47i3.75-0.47i 0.190.51i0.19-0.51i $0.40+0.55i$ 0.180.31i-0.18-0.31i 0.290.00i0.29-0.00i 1.871.33i1.87-1.33i 0.650.38i0.65-0.38i 0.50+0.20i-0.50+0.20i
4.861.69i4.86-1.69i $0.65+1.04i$ 0.620.21i0.62-0.21i 0.020.52i0.02-0.52i 0.230.00i0.23-0.00i $1.69+0.38i$ 0.770.57i0.77-0.57i 0.070.46i0.07-0.46i
7.511.15i7.51-1.15i 0.370.20i0.37-0.20i $1.71+1.96i$ 0.851.44i-0.85-1.44i 0.380.41i0.38-0.41i $0.86+1.19i$ $0.12+1.13i$ 0.31+0.55i-0.31+0.55i
8.290.78i8.29-0.78i 0.070.17i0.07-0.17i $1.12+1.75i$ 0.491.69i-0.49-1.69i 2.090.13i2.09-0.13i 1.70+0.08i-1.70+0.08i 0.63+0.04i-0.63+0.04i 0.330.12i0.33-0.12i
8.980.59i8.98-0.59i $0.02+0.06i$ 0.130.76i0.13-0.76i $0.04+0.60i$ 0.630.04i0.63-0.04i $0.06+0.67i$ 0.400.72i-0.40-0.72i 0.370.72i-0.37-0.72i
Ni ($n_{Y}=13$)
1.251.25i1.25-1.25i $0.17+0.03i$ 0.210.02i0.21-0.02i 0.090.01i0.09-0.01i 0.000.02i0.00-0.02i $1.39+1.40i$ 0.500.25i0.50-0.25i 0.070.54i-0.07-0.54i
4.552.09i4.55-2.09i $0.39+2.04i$ 1.213.38i1.21-3.38i 1.09+0.81i-1.09+0.81i 0.030.16i0.03-0.16i $18.07+14.32i$ 27.2329.25i-27.23-29.25i $11.13+14.58i$
4.900.62i4.90-0.62i 0.630.55i-0.63-0.55i 1.120.39i1.12-0.39i $0.31+0.14i$ $0.03+0.10i$ 0.210.77i-0.21-0.77i $1.00+1.50i$ 0.651.61i0.65-1.61i
6.662.47i6.66-2.47i $1.73+1.36i$ 1.540.52i-1.54-0.52i 0.280.48i0.28-0.48i 0.280.42i0.28-0.42i 1.716.13i-1.71-6.13i 4.76+13.35i-4.76+13.35i 4.636.51i4.63-6.51i
10.852.35i10.85-2.35i 0.51+1.33i-0.51+1.33i 1.251.49i1.25-1.49i 0.69+0.22i-0.69+0.22i 0.200.50i0.20-0.50i 10.734.15i-10.73-4.15i $13.80+11.53i$ 4.226.32i-4.22-6.32i
11.354.07i11.35-4.07i $0.66+1.06i$ 0.300.81i0.30-0.81i 0.56+0.08i-0.56+0.08i $0.56+0.48i$ 0.11+0.45i-0.11+0.45i 3.475.61i-3.47-5.61i $0.62+3.78i$
17.172.08i17.17-2.08i 0.16+0.10i-0.16+0.10i 0.300.07i0.30-0.07i 0.180.08i-0.18-0.08i 0.270.13i0.27-0.13i 2.46+1.61i-2.46+1.61i 0.485.19i0.48-5.19i 0.43+2.42i-0.43+2.42i
21.322.64i21.32-2.64i 0.070.04i0.07-0.04i 0.51+0.43i-0.51+0.43i 0.530.40i0.53-0.40i $1.00+0.80i$ $1.80+0.33i$ 4.363.54i-4.36-3.54i $0.62+2.95i$
23.063.11i23.06-3.11i 0.400.05i0.40-0.05i 0.74+0.24i-0.74+0.24i 0.300.23i0.30-0.23i $0.56+0.26i$ 2.15+1.85i-2.15+1.85i 6.044.15i6.04-4.15i 2.55+1.78i-2.55+1.78i
27.940.83i27.94-0.83i 0.260.46i-0.26-0.46i $0.29+0.92i$ 0.090.47i-0.09-0.47i $0.11+0.13i$ $0.85+0.94i$ 0.50+0.95i-0.50+0.95i 1.521.69i-1.52-1.69i
29.701.87i29.70-1.87i 0.08+0.05i-0.08+0.05i 0.050.21i-0.05-0.21i $0.21+0.10i$ $0.52+2.27i$ 2.340.51i-2.34-0.51i 2.582.94i2.58-2.94i 1.21+3.27i-1.21+3.27i
31.723.09i31.72-3.09i 0.040.08i-0.04-0.08i $0.16+0.15i$ 0.04+0.05i-0.04+0.05i $3.93+0.02i$ $0.33+2.39i$ 5.195.05i-5.19-5.05i $3.06+2.67i$
35.703.95i35.70-3.95i $0.11+0.06i$ 0.160.22i-0.16-0.22i $0.14+0.20i$ 2.540.62i2.54-0.62i $1.45+0.91i$ 5.540.51i-5.54-0.51i $2.42+0.27i$
Table 2: Parameters of the MPA($\mathbf{q}$}) representation of Eq. (15) fitted to the RPA data of Ca and Ni. These data generate the spectral $Y({\mathbf{q}},\omega)$ bands structures of Figs. 9 and $1$. Analogous MPA($\mathbf{q}$}) representations for the 26 studied metals are provided in Ref. [3].
In Fig. 11 we show the MPA($\mathbf{q}$}) spectral band structure of Ni, in this case modeled with 13 poles. All the parameters of the MPA($\mathbf{q}$}) representation can be found in Table 2. As for Ca, the most intense pole ($p=12$) is well defined, with the intensity spreading over a larger energy range. However, many poles, especially in the low-energy region (p=2p=2-66) around $10$ eV, shown in the zoomed panel of Fig. 6(d), exhibit residues that do not always comply with the relation $\mathrm{Re}[R_{p}(q)]>-\mathrm{Im}[R_{p}(q)]>0$. This could, in principle, be imposed in the fitting, as constraints on the residues $R_{p}$, similar to the ones imposed on the poles $\Omega_{p}$ through Eq. (16). However, such an approach sacrifices the accuracy of the MPA($\mathbf{q}$}) fit with respect to the numerical data, representing a considerable increase of the fitting error for Ni and many other metals. The fit with unconstrained residues allows for incoherent pole superposition and, therefore, underlines the complexity of plasmonic quasi-particle formation in non-homogeneous electronic materials like Ni.